# Atlas-free Brain Network Transformer

### v1

@@ Abstract
Current atlas-based approaches to brain network analysis rely heavily on standardized anatomical or connectivity-driven brain atlases. However, these fixed atlases often introduce significant limitations, such as spatial misalignment across individuals, functional heterogeneity within predefined regions, and atlas-selection biases, collectively undermining the reliability and interpretability of the derived brain networks. To address these challenges, we propose a novel atlas-free brain network transformer (atlas-free BNT) that leverages individualized brain parcellations derived directly from subject-specific resting-state fMRI data. Our approach computes ROI-to-voxel connectivity features in a standardized voxel-based feature space, which are subsequently processed using the BNT architecture to produce comparable subject-level embeddings. Experimental evaluations on sex classification and brain-connectome age prediction tasks demonstrate that our atlas-free BNT consistently outperforms state-of-the-art atlas-based methods, including elastic net, BrainGNN, Graphormer and the original BNT. Our atlas-free approach significantly improves the precision, robustness, and generalizability of brain network analyses. This advancement holds great potential to enhance neuroimaging biomarkers and clinical diagnostic tools for personalized precision medicine.
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI) has become an indispensable tool in neuroscience research, it enables the non-invasive mapping of brain function by tracking blood oxygen level-dependent (BOLD) signals over time. A critical step in many resting state-fMRI data analysis pipelines involves parcellating the brain into regions of interest (ROIs), which serve as the nodes for subsequent functional connectivity and brain network analyses. The standard workflow typically begins by spatially normalizing each subject’s brain images to a common coordinate system—most commonly the Montreal Neurological Institute (MNI) template—followed by segmenting the brain into distinct ROIs according to a “predefined” atlas. Such atlases are generally constructed using either prior anatomical knowledge or data-driven connectivity patterns among voxels. Anatomy-based atlases, such as the AAL-atlas [1] and Desikan-Killiany atlas [2], define regions based on structural landmarks such as gyri, sulci, or histological boundaries. In contrast, connectivity-based atlases, such as the Craddock atlas [3] and the Schaefer parcellation derived from the human connectome project [4], rely on fMRI time-series data to infer inter-voxel functional connectivities, grouping together voxels with similar temporal profiles through clustering techniques.
@@ 2.2 Deep Learning Approaches for Brain Network Analysis
Seeking to avoid the limitations of fixed brain atlases—such as spatial misalignment and intra-ROI heterogeneity—some studies have explored the direct analysis of preprocessed rs-fMRI volumes instead of relying on ROI-based functional connectivity. For example, Malkiel et al. [27] trained a multi-layer transformer to reconstruct fMRI data and fine-tuned the model for specific downstream tasks. Similarly, Sarraf et al. [28] applied a vision transformer to both rs-fMRI and structural MRI data for predicting Alzheimer’s disease progression. However, even after preprocessing, voxel-wise rs-fMRI data typically consists of over $\mathcal{O}(10^{5})$ noisy, spatially correlated time series per subject, posing significant computational challenges and increasing the risk of overfitting due to low signal-to-noise (SNR) ratio. Parcellating the brain into functionally coherent ROIs and averaging signals within each region greatly reduces data dimensionality and improves SNR. Moreover, analyzing functional connectivity between ROIs produces interpretable brain-network graphs that yield more meaningful insights into the brain’s functional organization.
@@ 3.1 Brain Parcellation via Clustering
    Spectral clustering partitions the voxels by leveraging the eigenspectrum of a similarity graph rather than relying on purely local, greedy merges. Starting from a similarity matrix $\bm{S}$}—the Pearson-correlation matrix—we construct a graph Laplacian ${\bm{L}}={\bm{D}}-{\bm{S}}$, where $\bm{D}$} is the diagonal degree matrix with $D_{ii}=\sum_{j}S_{ij}$. The top $k$ eigenvectors of $\bm{L}$} embed the voxels into a low-dimensional Euclidean space in which clusters become linearly separable. Sometimes the normalized variant of the Laplacian matrix ${\bm{L}}_{\textnormal{norm}}={\bm{D}}^{-\frac{1}{2}}{\bm{L}}{\bm{D}}^{-\frac{1}{2}}$ is preferred for a more balanced parcellation. The k-means clustering is then applied in this spectral space to obtain the final parcellation.
@@ 3.3 Atlas-free Brain Network Transformer
To extract standardized representations suitable for robust, cross-subject brain network analysis, we adopt the Brain Network Transformer (BNT) framework. BNT was originally developed to process atlas-based brain networks [23], thus it cannot directly handle the voxel-level representation in $\bm{Q}$}. To bridge this gap, we partition the 3D MNI space into overlapping blocks of size $K\times K\times K$ with a stride $s$, and treat each spatial block as an input node to BNT. The node feature vector ${\bm{x}}_{i}$ is then obtained by sum-pooling the voxel-wise features within the $i$-th block independently across all channels.
@@
Let ${\bm{X}}\in\mathbb{R}^{U\times V}$ represent the input node features, where $U$ is the the number of blocks/nodes and $V$ is the feature dimension. BNT employs a multi-head self-attention (MHSA) module consisting of $L$ layers to obtain attention-enhanced node embeddings. Specifically, the output from the $l$-th layer, denoted ${\bm{Z}}^{l}\in\mathbb{R}^{U\times V}$, is computed as follows:
@@
where ${\bm{Z}}_{m}^{l}$ is the output from the $m$-th attention head, $\{{\bm{W}}^{l,m}_{\mathcal{Q}},{\bm{W}}^{l,m}_{\mathcal{K}},{\bm{W}}^{l,m}_{\mathcal{V}},{\bm{W}}^{l}_{\mathcal{O}}\}$ are learnable model parameters, and $d^{l,m}_{\mathcal{K}}$ is the dimensionality of the key vectors for gradient stabilization. We can see that the query, key, and value matrices are all derived from the same input ${\bm{Z}}^{l-1}$, i.e. the output from the previous layer. After the MHSA modules, an orthonormal clustering readout function [23] aggregates node embeddings to generate a subject-level representation vector $\bm{h}$}. Finally, downstream brain network analysis tasks, such as classification or regression, are performed by feeding $\bm{h}$} into a multi-layer perceptron (MLP).
@@ 4.1 Experimental Setting
  • Emory Healthy Brain Study (EHBS). The EHBS is a longitudinal study focusing on healthy older adults aged 50–90, collecting multimodal neuroimaging data, and other health-related measures. Given that subjects’ biological ages change across follow-up sessions, multiple sessions from the same individual can be included in the brain-connectome age prediction dataset, provided all sessions from a given subject are consistently assigned to either the training or test set to prevent data leakage. Following quality control procedures, the dataset includes 2,255 sessions with recorded biological ages at the time of the MRI scan.
@@
We compare our proposed atlas-free method with four established atlas-based baseline approaches: (1) the classical elastic net with $l_{1}$ and $l_{2}$ norm regularizations [31]; (2) BrainGNN, which integrates ROI-selection pooling within a GNN architecture [22]; (3) Graphormer, which extends the transformer framework for graph representation learning [18] and (4) the original Brain Network Transformer (BNT), which leverages self-attention to learn connectivity patterns and the orthonormal clustering readout function to aggregate node embeddings [23]. Hyperparameters for these baseline methods are tuned and fully detailed in Appendix A. To mitigate atlas-selection bias inherent in traditional atlas-based analyses, we construct multiple brain networks using several widely adopted brain atlases. Specifically, we employ both anatomical and connectivity-driven atlases in the standard 2mm MNI space, including the anatomical AAL-atlas [1] and connectivity-based atlases such as Craddock-400 [3], Shen-368 [32], and the HCP multimodal parcellation with 360 areas (HCP-360) [33]. Fig. 5 illustrates the ROIs defined by those four atlases.
@@ 4.2 Sex Classification
Resting-state fMRI reveals intrinsic functional-connectivity patterns that systematically differ between males and females. This allows us to train a machine learning model to predict the biological sex from these connectivity features. The classification accuracy, sensitivity, specificity and the AUROC obtained from each approach are given in Table I. The proposed atlas-free BNT consistently outperforms all atlas-based baselines. Within the atlas-free framework, we observe that parcellations derived using agglomerative clustering (AC) yield slightly better performance than those produced by spectral clustering (SC). This advantage likely arises because AC is more effective at generating internally homogeneous clusters. However, AC operates in a stepwise, greedy fashion, merging two clusters at a time, and therefore requires considerably more computation time than SC. Across the atlas-based approaches, connectivity-driven atlases substantially outperform the anatomical AAL atlas, with the Craddock-400 atlas achieving the best overall results.
@@ 4.3 Brain Connectome Age Prediction
Figure 6: Atlas-free BNT’s predicted brain connectome age versus chronological age under two individualized brain parcellations: a) agglomerative clustering (AC) and b) spectral clustering (SC). Each point represents one subject/session. The mean and standard deviation of the absolute errors are calculated for both methods: a) $4.03\pm 2.99$ years; b) $4.06\pm 3.03$ years.
@@
Brain connectome age is an estimate of how old a person’s brain appears based on neurobiological connectivity features extracted from structural or functional neuroimaging data. It may differ from chronological age, reflecting individual variability in the brain’s aging process. The estimation model is typically trained on data from a cohort of healthy individuals, allowing brain connectome age to serve as a benchmark for normative brain development. A higher or lower brain connectome age relative to chronological age indicates accelerated or decelerated brain aging, respectively. This measure has been widely used as a biomarker to assess brain health and to aid in the early detection of neurodegenerative diseases. The mean absolute error (MAE) between the brain connectome age and chronological age is typically used as the evaluation metric for a prediction model. The obtained MAEs from different approaches are given in Table I. The atlas-free BNT consistently outperforms atlas-based methods, demonstrating greater robustness and generalizability. Figure 6 shows the atlas-free BNT’s predicted brain connectome age plotted against chronological age, and the figures for atlas-based approaches are presented in Appendix B. Among atlas-based approaches, connectivity-driven atlases generally achieve better performance than the anatomical AAL atlas, with the exception of BrainGNN. Across all atlas-based baselines, the Craddock-400 atlas remains the strongest performer.

### v2

@@ Abstract
Current atlas-based approaches to brain network analysis rely heavily on standardized anatomical or connectivity-driven brain atlases. However, these fixed atlases often introduce significant limitations, such as spatial misalignment across individuals, functional heterogeneity within predefined regions, and atlas-selection biases, collectively undermining the reliability and interpretability of the derived brain networks. To address these challenges, we propose a novel atlas-free brain network transformer (atlas-free BNT) that leverages individualized brain parcellations derived directly from subject-specific resting-state fMRI data. Our approach computes ROI-to-voxel connectivity features in a standardized voxel-based feature space, which are subsequently processed using the BNT architecture to produce comparable subject-level embeddings. Experimental evaluations on sex classification and brain-connectome age prediction tasks demonstrate that our atlas-free BNT consistently outperforms state-of-the-art atlas-based methods, including elastic net, BrainGNN, Graphormer and the original BNT. Our atlas-free approach significantly improves the precision, robustness, and generalizability of brain network analyses. This advancement holds great potential to enhance neuroimaging biomarkers and clinical diagnostic tools for personalized precision medicine. Reproducible code is available at https://github.com/shuai-huang/atlas_free_bnt
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI) has become an indispensable tool in neuroscience research, it enables the non-invasive mapping of brain function by tracking blood oxygen level-dependent (BOLD) signals over time. A critical step in many resting state-fMRI data analysis pipelines involves parcellating the brain into regions of interest (ROIs), which serve as the nodes for subsequent functional connectivity and brain network analyses. The standard workflow typically begins by spatially normalizing each subject’s brain images to a common coordinate system (most commonly the Montreal Neurological Institute (MNI) template) followed by segmenting the brain into distinct ROIs according to a “predefined” atlas. Such atlases are generally constructed using either prior anatomical knowledge or data-driven connectivity patterns among voxels. Anatomy-based atlases, such as the AAL-atlas [31] and Desikan-Killiany atlas [6], define regions based on structural landmarks such as gyri, sulci, or histological boundaries. In contrast, connectivity-based atlases, such as the Craddock atlas [3] and the Schaefer parcellation derived from the human connectome project [28], rely on fMRI time-series data to infer inter-voxel functional connectivities, grouping together voxels with similar temporal profiles through clustering techniques.
@@ 2.2 Deep Learning Approaches for Brain Network Analysis
Seeking to avoid the limitations of fixed brain atlases such as spatial misalignment and intra-ROI heterogeneity, some studies have explored the direct analysis of preprocessed rs-fMRI volumes instead of relying on ROI-based functional connectivity. For example, Malkiel et al. [23] trained a multi-layer transformer to reconstruct fMRI data and fine-tuned the model for specific downstream tasks. Similarly, Sarraf et al. [27] applied a vision transformer to both rs-fMRI and structural MRI data for predicting Alzheimer’s disease progression. However, even after preprocessing, voxel-wise rs-fMRI data typically consists of over $\mathcal{O}(10^{5})$ noisy, spatially correlated time series per subject, posing significant computational challenges and increasing the risk of overfitting due to low signal-to-noise (SNR) ratio. Parcellating the brain into functionally coherent ROIs and averaging signals within each region greatly reduces data dimensionality and improves SNR. Moreover, analyzing functional connectivity between ROIs produces interpretable brain-network graphs that yield more meaningful insights into the brain’s functional organization.
@@ 3.1 Brain Parcellation via Clustering
    Spectral clustering partitions the voxels by leveraging the eigenspectrum of a similarity graph rather than relying on purely local, greedy merges. Starting from a similarity matrix $\bm{S}$}, i.e. the Pearson-correlation matrix, we construct a graph Laplacian ${\bm{L}}={\bm{D}}-{\bm{S}}$, where $\bm{D}$} is the diagonal degree matrix with $D_{ii}=\sum_{j}S_{ij}$. The top $k$ eigenvectors of $\bm{L}$} embed the voxels into a low-dimensional Euclidean space in which clusters become linearly separable. In this paper we use the normalized variant of the Laplacian matrix ${\bm{L}}_{\textnormal{norm}}={\bm{D}}^{-\frac{1}{2}}{\bm{L}}{\bm{D}}^{-\frac{1}{2}}$ for a more balanced parcellation. The k-means clustering is then applied on the top $k$ eigenvectors of ${\bm{L}}_{\textnormal{norm}}$ in this spectral space to obtain the final parcellation.
@@ 3.3 Atlas-free Brain Network Transformer
To extract standardized representations suitable for robust, cross-subject brain network analysis, we adopt the Brain Network Transformer (BNT) framework. BNT was originally developed to process atlas-based brain networks [14], thus it cannot directly handle the voxel-level representation in $\bm{Q}$}. To bridge this gap, we partition the 3D MNI space into overlapping blocks of size $B\times B\times B$ with a stride $s$, and treat each spatial block as an input node to BNT. The node feature vector ${\bm{x}}_{i}$ is then obtained by sum-pooling the voxel-wise features within the $i$-th block independently across all channels.
@@
Let ${\bm{X}}\in\mathbb{R}^{U\times V}$ represent the input node features, where $U$ is the the number of blocks/nodes and $V$ is the feature dimension. BNT employs a multi-head self-attention (MHSA) module consisting of $\mathcal{L}$ layers to obtain attention-enhanced node embeddings. Specifically, the output from the $l$-th layer, denoted ${\bm{Z}}^{l}\in\mathbb{R}^{U\times V}$, is computed as follows:
@@
where ${\bm{Z}}_{m}^{l}$ is the output from the $m$-th attention head, $\{{\bm{W}}^{l,m}_{\mathcal{Q}},{\bm{W}}^{l,m}_{\mathcal{K}},{\bm{W}}^{l,m}_{\mathcal{V}},{\bm{W}}^{l}_{\mathcal{O}}\}$ are learnable model parameters, and $d^{l,m}_{\mathcal{K}}$ is the dimensionality of the key vectors for gradient stabilization. We can see that the query, key, and value matrices are all derived from the same input ${\bm{Z}}^{l-1}$, i.e. the output from the previous layer.
After the MHSA encoder, we apply the orthonormal clustering readout (OCR) proposed in [14] to aggregate the final-layer node embeddings $\bm{Z}$$^{\mathcal{L}}$ into a compact, subject-level representation $\bm{H}$}. In OCR, the node embeddings are softly clustered into $K$ groups with learnable centroids ${\bm{E}}\in\mathbb{R}^{K\times V}$. The soft assignment of node $i$ to cluster $k$ is computed via a softmax over inner products,
${\bm{P}}_{i,k}=\frac{\exp\left(\langle{\bm{Z}}_{i}^{L},{\bm{E}}_{k}\rangle\right)}{\sum_{k^{\prime}=1}^{K}\exp\left(\langle{\bm{Z}}_{i}^{L},{\bm{E}}_{k^{\prime}}\rangle\right)},$ (8)
where $\langle{\bm{Z}}_{i}^{\mathcal{L}},{\bm{E}}_{k}\rangle$ denotes the inner product between the $i$-th node embedding and the $k$-th centroid. Given ${\bm{P}}\in\mathbb{R}^{U\times K}$, the graph-level embedding is obtained by pooling node features within each cluster:
${\bm{H}}={\bm{P}}^{T}{\bm{Z}}$ (9)
Finally, we vectorize ${\bm{H}}\in\mathbb{R}^{K\times V}$ (i.e., ${\bm{h}}=\textnormal{flatten}({\bm{H}})$) and feed it into a multi-layer perceptron (MLP) for downstream tasks such as classification or regression.
@@ 4.1 Experimental Setting
  • Emory Healthy Brain Study (EHBS). The EHBS is a longitudinal study focusing on healthy older adults aged 50–90. Multimodal neuroimaging data and other health-related measures are collected, after obtaining written consent from the subjects and receiving approval from the Institutional Review Board (IRB) of Emory University. Given that subjects’ biological ages change across follow-up sessions, multiple sessions from the same individual can be included in the brain-connectome age prediction dataset, provided all sessions from a given subject are consistently assigned to either the training or test set to prevent data leakage. Following quality control procedures, the dataset includes 2,255 sessions with recorded biological ages at the time of the MRI scan.
@@
We compare our proposed atlas-free method with four established atlas-based baseline approaches: (1) the classical elastic net with $l_{1}$ and $l_{2}$ norm regularizations [39]; (2) BrainGNN, which integrates ROI-selection pooling within a GNN architecture [22]; (3) Graphormer, which extends the transformer framework for graph representation learning [38] and (4) the original Brain Network Transformer (BNT), which leverages self-attention to learn connectivity patterns and the orthonormal clustering readout (OCR) function to aggregate node embeddings [14]. Hyperparameters for these baseline methods are tuned and fully detailed in Appendix A. To mitigate atlas-selection bias inherent in traditional atlas-based analyses, we construct multiple brain networks using several widely adopted brain atlases. Specifically, we employ both anatomical and connectivity-driven atlases in the standard 2mm MNI space, including the anatomical AAL-atlas [31] and connectivity-based atlases such as Craddock-400 [3], Shen-368 [30], and the HCP multimodal parcellation with 360 areas (HCP-360) [10]. Fig. 5 illustrates the ROIs defined by those four atlases.
@@ 4.2 Sex Classification
Resting-state fMRI reveals intrinsic functional-connectivity patterns that systematically differ between males and females. This allows us to train a machine learning model to predict the biological sex from these connectivity features. For this classification task, elastic net adopts the logistic regression formulation for training, all other classification models were trained by minimizing the cross-entropy loss. The classification accuracy, sensitivity, specificity and the AUROC obtained from each approach are given in Table I. The proposed atlas-free BNT consistently outperforms all atlas-based baselines. Within the atlas-free framework, we observe that parcellations derived using agglomerative clustering (AC) yield slightly better performance than those produced by spectral clustering (SC). This advantage likely arises because AC is more effective at generating internally homogeneous clusters. However, AC operates in a stepwise, greedy fashion, merging two clusters at a time, and therefore requires considerably more computation time than SC. Across the atlas-based approaches, connectivity-driven atlases substantially outperform the anatomical AAL atlas, with the Craddock-400 atlas achieving the best overall results.
@@ 4.3 Brain Connectome Age Prediction
Figure 6: Atlas-free BNT’s predicted brain-connectome age versus chronological age under two individualized brain parcellations: a) agglomerative clustering (AC) and b) spectral clustering (SC). Each point represents one subject/session. The mean and standard deviation of the absolute errors are calculated for both methods: a) $4.03\pm 2.99$ years; b) $4.06\pm 3.03$ years.
@@
Brain-connectome age is an estimate of how old a person’s brain appears based on neurobiological connectivity features extracted from structural or functional neuroimaging data. It may differ from chronological age, reflecting individual variability in the brain’s aging process. The estimation model is typically trained on data from a cohort of healthy individuals, allowing brain-connectome age to serve as a benchmark for normative brain development. A higher or lower brain-connectome age relative to chronological age indicates accelerated or decelerated brain aging, respectively. This measure has been widely used as a biomarker to assess brain health and to aid in the early detection of neurodegenerative diseases. For this regression task, elastic net follows a standard linear regression formulation, all other regression models were trained by minimizing the mean squared error (MSE) between the brain-connectome age and chronological age. For evaluation, we report mean absolute error (MAE), which is commonly used in brain-age studies because it provides a more intuitive measure of prediction error in years than MSE. The obtained MAEs from different approaches are also given in Table I. The atlas-free BNT consistently outperforms atlas-based methods, demonstrating greater robustness and generalizability. Figure 6 shows the atlas-free BNT’s predicted brain-connectome age plotted against chronological age, and the figures for atlas-based approaches are presented in Appendix B. Among atlas-based approaches, connectivity-driven atlases generally achieve better performance than the anatomical AAL atlas, with the exception of BrainGNN. Across all atlas-based baselines, the Craddock-400 atlas remains the strongest performer.
4.4 Ablation Study
Table II: Ablation study for the proposed atlas-free BNT (AC).
BNT Sex classification Brain age prediction
Method FNN MHSA OCR Accuracy Sensitivity Specificity AUROC Mean Absolute Error
[…]
We conduct an ablation study to assess how individual components of the atlas-free BNT contribute to predictive performance. Specifically, we choose agglomerative clustering (AC) for brain parcellation and evaluate the following model variants, each obtained by removing or simplifying a single module while keeping the remaining pipeline unchanged:
  • V1 (No FNN feature transform): In the full model, the feedforward neural network (FNN) maps the connectivity feature vector $\bm{g}$} to a latent representation $\bm{q}$} that is better suited for subsequent block-wise node feature construction. In this variant, we remove the FNN and form node features by directly sum-pooling $\bm{g}$} within each spatial block.
  • V2 (No BNT; MLP on pooled tokens): The BNT module comprises multi-head self-attention (MHSA) and the orthonormal clustering readout (OCR), it models interactions among block tokens and aggregates them into a subject-level embedding. In this variant, we remove the entire BNT module and instead concatenate all block-level node features and feed the resulting vector directly into the downstream MLP regressor.
  • V3 (No OCR readout): In the full model, OCR summarizes MHSA-updated node features via soft clustering to produce the final graph-level embedding. In this variant, we remove OCR and replace it with a standard MHSA-only aggregation for prediction (i.e., the transformer encoder is retained, but the OCR-based readout is omitted).
  • V4 (No MHSA transformer): In the full model, the MHSA transformer captures the long-range dependencies and inter-block interactions, it infuses global information into the node representations before they are clustered by OCR. In this variant, we remove the MHSA encoder and apply OCR directly to the original block-level node features for downstream prediction.
Across both tasks, the ablation results in Table II suggest that different components contribute in complementary ways, with the overall framework being most beneficial when all modules are used together. For sex classification, performance is relatively stable across the four variants, and the full atlas-free BNT remains the top performer. This pattern indicates that sex-related features are likely strong and broadly distributed in the connectivity representation, so even simplified aggregation strategies can recover much of the discriminative information; nevertheless, retaining the full pipeline provides a consistent edge, likely by better standardizing features and reducing sensitivity to subject-specific parcellation variability.
In contrast, brain-age prediction shows clear separation among the four variants, implying that age-related effects are more subtle and depend more heavily on how information is integrated across the brain. In particular, the full model achieves an MAE of ($4.06\pm 0.21$) years. Removing the BNT module entirely and feeding concatenated block features directly to an MLP (V2) yields the largest performance drop ($4.95\pm 0.27$). This shows that modeling inter-block relationships and performing structured global aggregation via BNT are critical for accurate age prediction. In contrast, removing the OCR readout while retaining MHSA (V3) produces essentially no degradation ($4.05\pm 0.21$). It suggests that for this task the transformer encoder alone already learns a sufficiently informative global representation and that OCR provides only marginal additional benefit. Eliminating the MHSA encoder and applying OCR directly to the uncontextualized block features (V4) increases error to $4.20\pm 0.22$, which demonstrates the key role of self-attention in encoding each block embedding with whole-brain context before pooling. Finally, bypassing the FNN feature transform and pooling the original connectivity vectors within blocks (V1) also degrades performance ($4.28\pm 0.20$), indicating that the learned non-linear mapping helps produce block-level features that are more suitable for subsequent feature aggregation. Overall, these results reinforce that the atlas-free BNT’s main advantage (especially for brain-age prediction) comes from transformer-based global context modeling, while the FNN and OCR contribute additional gains that are smaller but complementary.